In a recent​ year, person A and person B had two concert​ tours, and together they generated $197.6 million in ticket sales. If person B took in $18.6
million less than person​ A, how many millions of dollars did each tour​ generate?

Answer by Timnewman(323)

From your first sentence,
A+B=197.6-----(1)
From your second sentence,
B=A-18.6------(2)
put 2 in 1
A+A-18.6=197.6
2A=197.6+18.6
2A=216.2
A=108.1
Put A=108.1 in 2
B=108.1-18.6
B=89.5
Person A generated $108.1
Person B generated $89.5
HOPE THIS HELPS?